# Easy attention: A simple attention mechanism for temporal predictions with transformers

### v3

@@ Introduction
where $\hat{\mathbf{x}}=(\hat{x},\hat{y},\hat{z})$ and $\mathbf{x}=(x,y,z)$ are the model prediction and the ground truth from numerical solver, respectively. Following the assessment in Ref. [30], we focus on the short-time accuracy and compute the error on the first 512 time steps of the test sequence. The time window in which the error is computed is then $T_{\varepsilon}=5.12$. We also assess the computation time of training ($t_{c}$) and number of floating-point operations ($N_{f}$) for one forward propagation with batch size of 1. We summarize the results in the supplementary material, where the proposed model using dense easy attention yields the lowest error, 1.99%, demonstrating the superior performance of the easy-attention method applied on temporal-dynamics predictions. The LSTM model has the highest number of learnable parameters, but the lowest training time and number of floating-point operations. However, it yields the highest $l_{2}$-norm error of 37.68%, which indicates that the LSTM model is not able to reproduce the correct dynamical behaviour over a long time given the same set up as the transfomer. The dense-easy-attention model reduces the time for training by 17% and the computational complexity by 25% with respect to the self-attention-based model while surpassing the number of learnable parameters of the self-attention-based model. This is due to the fact that the dense-easy-attention module comprises $\bm{\alpha}\in\mathbb{R}^{h\times p\times p}$ and $\mathbf{W}_{V}\in\mathbb{R}^{p\times d_{\rm model}}$ while the self-attention module comprises $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V},\mathbf{W}_{C}\in\mathbb{R}^{p\times d_{\rm model}}$ which has fewer parameters in the present study. However the computational cost for the easy-attention models is lower than that of the self attention since the softmax and the inner product inside the latter are eliminated. Note that the model employing sparse easy attention exhibits an error of 2.79% while only having 52% of the learnable parameters of the self-attention-based transformer, which implies that the sparsification induced by an offset of 0 benefits the robustness of the easy-attention method applied to temporal-dynamics prediction. The sparsity is defined as follows: the $\bm{\alpha}$ matrix has a number of non-zero diagonals equal to $12\cdot\rm{offset}+$ 1. We also implement the averaged relative Euclidean norm error $\psi(t)$ between the ground truth and the prediction, which is proposed in Ref. [43]:
@@
where $\langle\cdot\rangle_{\rm ens}$ and $\langle\cdot\rangle_{t}$ indicate ensemble averaging over 100 sets of time series and over 10,000 time steps, respectively. Note that the set of time series used for evaluation are again generated by integrating the system with new initial states of $x_{0}=y_{0}=z_{0}=6$ with independent perturbations $\epsilon_{x}$, $\epsilon_{y}$ and $\epsilon_{z}$ where $\epsilon_{i}\sim\mathcal{N}(0,1)$ for each variable. Fig. 2 depicts the evolution $\psi(t)$ for the different models. One can observe that the easy-attention-based transformer deviates from the ground truth later than the LSTM and the self-attention-based transformer in terms of short-term prediction. We adopt a threshold of $\psi$ = 0.4 as an adequate agreement with respect to the reference trajectory, it can be observed that the transformers using easy attention and sparse easy attention produce accurate instantaneous predictions for 7.04 and 5.97 time units, respectively. However, the self-attention-based transformer and the LSTM yields 4.90 and 1.10 accurate instantaneous predictions, respectively. The results indicate that the easy-attention-based transformer outperforms the self-attention-based transformer and the LSTM for short-term prediction.
Figure 2: Assessment of short-term predictions for the various models. Relative Euclidean norm of error $\psi(t)$ averaged over 100 sets of time series comprising 10,000 time steps, generated by integrating the system with new initial states of $x_{0}=y_{0}=z_{0}=6$ with independent perturbations $\epsilon_{x}$, $\epsilon_{y}$ and $\epsilon_{z}$ where $\epsilon_{i}\sim\mathcal{N}(0,1)$ for each variable. The dashed horizontal line shows the threshold value considered for accurate predictions, namely $\epsilon$ = 0.4.
@@
Figure 3: Characterization of the dynamics of the Lorenz system.(Top) The Lorenz map representing the two consecutive local maxima of $z$ produced by each model. Note that we amplify the region where $\max_{n}(z)\in(37.3,38.3)$ and $\max_{n+1}(z)\in(40,43)$. (Bottom) Leading Lyapunov exponents obtained by adding a perturbation at time step of 400 with a magnitude of $\delta=1\times 10^{-6}$. Note that we report the obtained Lyapunov exponents by ensemble averaging 10 perturbed test dataset with initial states of $x_{0}=y_{0}=z_{0}=6$ whereas ODE denotes the results obtained by Lorenz system. The obtained Lyapunov exponents were scaled by the adopted time step $\Delta t=0.01$ in the present study, and $\Delta t$ denotes the time after the introduction of the perturbation. The dashed lines denote the fitted exponential evolution used to obtain the corresponding Lyapunov exponents.
Furthermore, Fig. 4 depicts the probability density functions (PDF) of the $x$ and $z$ variables. To highlight the entire path and avoid focusing near $(0,0,0)$, an arc-length parametrization is performed to ${\bf x}(t)=\big{(}x(t),y(t),z(t)\big{)}$ to obtain a natural curve ${\bf\hat{x}}(s)={\bf x}[t(s)]$. This is done by considering $t(s)$, the inverse function of $s(t)=\int_{0}^{t}\|{\bf x}^{\prime}(r)\|_{2}{\rm dr}$. As before, the models with the most similar PDF to the one of the test data are the two easy-attention-based transformers. Both exhibit a better performance when considering the symmetry on the $x$ axis, as shown below. It is necessary to mention the importance of replicating symmetries when learning chaotic systems as the main exoticism of complex systems relies on their unique structures, spatially; attractors and spatio-temporally, probability density functions. Being able to reproduce such structures it is not a trivial consequence of the training procedure: while the model has as single objective temporal predictions, inspecting the results in Fig. 4 one can conclude that not only the space is well reconstructed but the velocity, the time derivative of the state variable too, as the frequency per volume unit is also learned. In Fig. 4 we depict the embedded manifold given by the joint PDF (JPDF). Note the excellent agreement with the ground truth observed after the symmetry transformation. The possibility of accurately reproducing spatio-temporal structures will become critical in more challenging chaotic problems such as turbulence.
@@
Compared with the self-attention-based transformer, the transformers using easy attention have the lowest computational cost for training and exhibit higher accuracy on reproduction of long-term dynamics, which indicates that the transformer models using easy attention are more robust for temporal-dynamics prediction than the model using self attention.
Low-order model of turbulence shear flows
@@
As introduced above, we will investigate the transformer’s ability to reconstruct and capture both temporal and spatial dynamics for this turbulent shear flow. We will use essentially the same architecture as that used for the Lorenz system, changing the embedding module from time2vec to the time-space embedding module proposed in Refs. [31, 45], see the Supplementary Material for additional details. We will also compare these results with the one-layer LSTM architecture reported in Ref. [43].
@@
Figure 5: Turbulent-flow predictions by the various methods. (Top) Probability density function of the Poincaré maps showing the intersection with $a_{2}=0$ plane (with ${\rm d}a_{2}/{\rm d}t<0)$. In all the cases black denotes the ground truth while the colors correspond to predictions based on (from left to right): eay attention, sparse easy attention, self attention and LSTM. (Bottom) Turbulence statistics corresponding to (left) streamwise mean profile, (middle) streamwise velocity fluctuations and (right) Reynolds shear stress. Colors correspond to the cases on the top panels.
@@
To conclude the assessment of our easy-attention-based model, we will consider a more complex problem of industrial relevance: a model of a nuclear reactor. In particular, we will evaluate the feasibility of using easy attention as a surrogate model to predict the temporal evolution of the state variables in the reactor in order to increase safety and decrease operational costs. The pebble-bed fluoride-salt-cooled high-temperature (PB-FHR) reactor under study here is a Generation-IV nuclear reactor [47] that utilizes a liquid salt coolant, solid fuel pebbles and a graphite moderator. The fluoride-salt coolant has attractive properties that motivate the FHR design both economically and safety-wise, including high boiling point, low operational pressure and high heat-transfer efficiency [48]. We will consider the generic FHR (gFHR), a publicly accessible benchmark model created by Kairos Power [41] which retains the main neutronics and thermal-hydraulics information of their proprietary Karios Power FHR (KP-FHR) model. Data is generated with the System Analysis Module (SAM), a tool developed by Argonne National Laboratory for whole-plant transient analysis of advanced reactor concepts [49]. The SAM gFHR model considered here was developed using a hybrid method [50] that combines high-fidelity neutronics simulations and intermediate-fidelity thermal-hydraulics simulations for an entire-plant representation of the reactor configuration. In this study we will focus on predicting the temporal evolution of 13 state variables, and more details regarding the employed datasets can be found in the Supplementary Material.
@@
Summary and conclusions
We proposed a novel attention mechanism for transformer neural networks, namely the easy attention, which does not rely on queries and keys to generate attention scores and eliminates the nonlinear $\rm softmax$ function. The easy attention arises from the nonelastic nature of the trace, as this makes impossible any backwards traceability of the information, decreasing the relevance of the exact value of each matrix element. Taking this into account, we propose to learn the attention scores $\bm{\alpha}$ directly, instead of learning the query and key matrices, removing the $\mathbf{W}_{K}$, $\mathbf{W}_{Q}$ and $\rm softmax$ from the attention mechanism, yielding a considerable reduction in the numbers of parameters. Furthermore, we propose the sparse-easy-attention method, which reduces the number of learnable parameters of the tensor $\bm{\alpha}$ by only learning the (near-) diagonal elements. We also propose a method to perform multi-head easy attention which employs multiple $\bm{\alpha}$ to improve the long-term predictions. Considering chaotic systems as our predictive problem, the invariant nature of attractors and the sensitivity to small perturbations on the initial conditions, suggest that it may be more convenient to use an input-independent model. In future works we will expand the interpretability of the model with operator theory [51], leveraging ideas from both Koopman Operator Theory and quantum mechanics, specifically the Schrödinger equation, where the operator evolving the observable is input independent.
@@
Subsequently, we apply easy attention in the proposed transformer model for temporal-dynamics prediction of chaotic systems and we investigate the performance of our models on the Lorenz system. The transformer model with dense easy attention yields the lowest relative $l_{2}$-norm error of 2.0%, outperforming the same transformer architecture with the self attention (7.4%) and the LSTM model (37.7%). Furthermore, when inspecting the nine-equation model of wall-bounded turbulence by Moehlis et al. [40] the long-term statistics discussed in the Supplementary Material show how LSTM yields very similar performance in the mean flow compared with the easy attention. Regarding the PDF of the Poincaré section shown in Fig. 5, the easy attention also exhibits better predictions than the other models, in particular when accounting for the existing symmetries observed in Fig. 4 (top). Therefore, the easy attention arises as the most accurate model when studying chaotic systems. Regarding the studied model of a nuclear reactor, the easy attention shows the best reconstructive capability when considering PSD and PDF analysis. The results also show that, compared with the self-attention-based transformer, the transformers using easy attention have a much lower computational cost, exhibit less complexity and have less parameters. Thus the transformer model using easy attention is more robust for temporal-dynamics predictions.
Furthermore, a significant advantage of the easy-attention-based transformer is its accurate reconstruction of the invariant sets, as shown in Figs. 4,6. These invariant sets arise when inspecting the systems independently of time, as the latter converge to a stationary probability density function and a fixed attractor in state space. As the PDF is reconstructed, although our temporal prediction diverges, the density of information per neighborhood is accurately represented by the easy-attention model. Taking into account the shadowing lemma [52], it can be stated that the prediction given by the transformer can be trusted to represent the orbits of the dynamical system, since the diverged orbit remains as a pseudo-orbit infinitely close to another orbit on the original system.
@@
Methods
In this section, we will focus on the mathematical development of the attention part of the model and its implications on the final output, since the other blocks follow the standard transformer architecture [26]. To this end we will first examine the self-attention model an later explain in detail the easy-attention method. The mathematical implementation of a transformer encoder block is expressed as follows [53]: a transformer block is a parameterised function $T:\mathbb{R}^{n\times d}\rightarrow\mathbb{R}^{n\times d}$, usually consisting of an attention part and a feed forward-network, both with residual connections and normalization layer, see the transformer block used for each case in the Supplementary Material.

### v4

@@ III Temporal-dynamics predictions of chaotic systems using a transformer
where $\hat{\mathbf{x}}=(\hat{x},\hat{y},\hat{z})$ and $\mathbf{x}=(x,y,z)$ are the model prediction and the ground truth from numerical solver, respectively. Following the assessment in Ref. [32], we focus on the short-time accuracy and compute the error on the first 512 time steps of the test sequence. The time window in which the error is computed is then $T_{\varepsilon}=5.12$ convective time units.We also assess the computation time of training ($t_{c}$) and number of floating-point operations ($N_{f}$) for one forward propagation with batch size of 1. We summarize the results in the supplementary material in Tab. 5, where the proposed model using dense easy attention yields the lowest error, 1.99%, demonstrating the superior performance of the easy-attention method applied on temporal-dynamics predictions. The LSTM model has the highest number of learnable parameters, but the lowest training time and number of floating-point operations. However, it yields the highest $l_{2}$-norm error of 37.68%, which indicates that the LSTM model is not able to reproduce the correct dynamical behaviour over a long time given the same set up as the transfomer. The dense-easy-attention model reduces the time for training by 17% and the computational complexity by 25% with respect to the self-attention-based model while surpassing the number of learnable parameters of the self-attention-based model. This is due to the fact that the dense-easy-attention module comprises $\bm{\alpha}\in\mathbb{R}^{h\times p\times p}$, where $h$ is the number of heads, $p$ the size of the input after embedding, and $\mathbf{W}_{V}\in\mathbb{R}^{p\times d_{\rm model}}$ while the self-attention module comprises $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}\in\mathbb{R}^{p\times d_{\rm model}}$ which has fewer parameters in the present study. However the computational cost for the easy-attention models is lower than that of the self attention since the softmax and the inner product inside the latter are eliminated. Note that the model employing sparse easy attention exhibits an error of 2.79% while only having 52% of the learnable parameters of the self-attention-based transformer, which implies that the sparsification induced by an offset of 0 benefits the robustness of the easy-attention method applied to temporal-dynamics prediction. The sparsity is defined as follows: the $\bm{\alpha}$ matrix has a number of non-zero diagonals equal to $12\cdot\rm{offset}+$ 1. We also implement the averaged relative Euclidean norm error $\psi(t)$ between the ground truth and the prediction, which is proposed in Ref. [17]:
@@
where $\langle\cdot\rangle_{\rm ens}$ and $\langle\cdot\rangle_{t}$ indicate ensemble averaging over 100 sets of time series and over 10,000 time steps, respectively. Note that the set of time series used for evaluation are again generated by integrating the system with new initial states of $x_{0}=y_{0}=z_{0}=6$ with independent perturbations $\epsilon_{x}$, $\epsilon_{y}$ and $\epsilon_{z}$ where $\epsilon_{i}\sim\mathcal{N}(0,1)$ for each variable. Fig. 2 depicts the evolution $\psi(t)$ for the different models. One can observe that the easy-attention-based transformer deviates from the ground truth later than the LSTM and the self-attention-based transformer in terms of short-term prediction. We adopt a threshold of $\psi$ = 0.4 as an adequate agreement with respect to the reference trajectory, it can be observed that the transformers using easy attention and sparse easy attention produce accurate instantaneous predictions for 7.04 and 5.97 time units, respectively. However, the self-attention-based transformer and the LSTM yields 4.90 and 1.10 respectively. The results indicate that the easy-attention-based transformer outperforms the self-attention-based transformer and the LSTM for short-term prediction.
Figure 2: Assessment of short-term predictions for the various models. Relative Euclidean norm of error $\psi(t)$ averaged over 100 sets of time series comprising 10,000 time steps, generated by integrating the system with new initial states of $x_{0}=y_{0}=z_{0}=6$ with independent perturbations $\epsilon_{x}$, $\epsilon_{y}$ and $\epsilon_{z}$ where $\epsilon_{i}\sim\mathcal{N}(0,1)$ for each variable. The dashed horizontal line shows the threshold value considered for accurate predictions, namely $\psi(t)$ = 0.4.
@@
Figure 3: Characterization of the dynamics of the Lorenz system.(Top) The Lorenz map representing the two consecutive local maxima of $z$ produced by each model. Note that we amplify the region where $\max_{n}(z)\in(37.3,38.3)$ and $\max_{n+1}(z)\in(40,43)$. (Bottom) Leading Lyapunov exponents obtained by adding a perturbation at time step of 400 with a magnitude of $\delta=1\times 10^{-6}$. Note that we report the obtained Lyapunov exponents by ensemble averaging 10 perturbed test dataset with initial states of $x_{0}=y_{0}=z_{0}=6$ whereas ODE denotes the results obtained by Lorenz system. The obtained Lyapunov exponents were scaled by the adopted time step $\Delta t=0.01s$ in the present study, and $\Delta t$ denotes the time after the introduction of the perturbation. The dashed lines denote the fitted exponential evolution used to obtain the corresponding Lyapunov exponents.
Furthermore, Fig. 4 depicts the probability density functions (PDF) of the $x$ and $z$ variables. To highlight the entire path and avoid focusing near $(0,0,0)$, an arc-length parametrization is performed to ${\bf x}(t)=\big{(}x(t),y(t),z(t)\big{)}$ to obtain a natural curve ${\bf\hat{x}}(s)={\bf x}[t(s)]$. This is done by considering $t(s)$, the inverse function of $s(t)=\int_{0}^{t}\|{\bf x}^{\prime}(r)\|_{2}{\rm dr}$. As before, the models with the most similar PDF to the one of the test data are the two easy-attention-based transformers. Both exhibit a better performance when considering the symmetry on the $x$ axis, as shown below. It is necessary to mention the importance of replicating symmetries when learning chaotic systems as the main exoticism of complex systems relies on their unique structures or invariant-sets, spatio-temporally; attractors and probability density functions. Being able to reproduce such structures it is not a trivial consequence of the training procedure: while the model has as single objective temporal predictions, inspecting the results in Fig. 4 one can conclude that not only the space is well reconstructed but the velocity, the time derivative of the state variable too, as the frequency per volume unit is also learned. In Fig. 4 we depict the embedded manifold given by the joint PDF (JPDF). Note the excellent agreement with the ground truth observed after the symmetry transformation. The possibility of accurately reproducing spatio-temporal structures will become critical in more challenging chaotic problems such as turbulence.
@@
Compared with the self-attention-based transformer, the transformers using easy attention have lower computational cost for training and exhibit higher accuracy on reproduction of long-term dynamics, which indicates that the transformer models using easy attention are more robust for temporal-dynamics prediction than the model using self attention.
Turbulence shear flows
@@
As introduced above, we will investigate the transformer’s ability to reconstruct and capture both temporal and spatial dynamics for this turbulent shear flow. We will use essentially the same architecture as that used for the Lorenz system, changing the embedding module from time2vec to the time-space embedding module proposed in Refs. [33, 46], see Tab. 6 in the Supplementary Material for additional details. We will also compare these results with the one-layer LSTM architecture reported in Ref. [17].
@@
Figure 5: Turbulent-flow predictions by the various methods. (Top) Probability density function of the Poincaré maps showing the intersection with $a_{2}=0$ plane (with ${\rm d}a_{2}/{\rm d}t<0)$. In all the cases black denotes the ground truth while the colors correspond to predictions based on (from left to right): easy attention, sparse easy attention, self attention and LSTM. (Bottom) Turbulence statistics corresponding to (left) streamwise mean profile, (middle) streamwise velocity fluctuations and (right) Reynolds shear stress. Colors correspond to the cases on the top panels.
@@
To conclude the assessment of our easy-attention-based model, we will consider a more complex problem of industrial relevance: a model of a nuclear reactor. In particular, we will evaluate the feasibility of using easy attention as a surrogate model to predict the temporal evolution of the state variables in the reactor in order to increase safety and decrease operational costs. The pebble-bed fluoride-salt-cooled high-temperature (PB-FHR) reactor under study here is a Generation-IV nuclear reactor [48] that utilizes a liquid salt coolant, solid fuel pebbles and a graphite moderator. The fluoride-salt coolant has attractive properties that motivate the FHR design both economically and safety-wise, including high boiling point, low operational pressure and high heat-transfer efficiency [49]. We will consider the generic FHR (gFHR), a publicly accessible benchmark model created by Kairos Power [43] which retains the main neutronics and thermal-hydraulics information of their proprietary Karios Power FHR (KP-FHR) model. Data is generated with the System Analysis Module (SAM), a tool developed by Argonne National Laboratory for whole-plant transient analysis of advanced reactor concepts [50]. The SAM gFHR model considered here was developed using a hybrid method [51] that combines high-fidelity neutronics simulations and intermediate-fidelity thermal-hydraulics simulations for an entire-plant representation of the reactor configuration. In this study we will focus on predicting the temporal evolution of 13 state variables, and more details regarding the employed datasets can be found in the Supplementary Material in Tabs. 8 & 9
@@
IV Summary and conclusions
We proposed a novel attention mechanism for transformer neural networks, namely the easy attention, which does not rely on queries and keys to generate attention scores and eliminates the nonlinear $\rm softmax$ function. The easy attention arises from the nonelastic nature of the trace, as this makes impossible any backwards traceability of the information, decreasing the relevance of the exact value of each matrix element. Taking this into account, we propose to learn the attention scores $\bm{\alpha}$ directly, instead of learning the query and key matrices, removing the $\mathbf{W}_{K}$, $\mathbf{W}_{Q}$ and $\rm softmax$ from the attention mechanism, yielding a considerable reduction in the numbers of parameters. Furthermore, we propose the sparse-easy-attention method, which further reduces the number of learnable parameters of the tensor $\bm{\alpha}$ by only learning the (near-) diagonal elements. We also propose a method to perform multi-head easy attention which employs multiple $\bm{\alpha}$ to improve the long-term predictions. Considering chaotic systems as our predictive problem, the invariant nature of attractors and the sensitivity to small perturbations on the initial conditions, suggest that it may be more convenient to use an input-independent model. In future works we will expand the interpretability of the model with operator theory [52], leveraging ideas from both Koopman Operator Theory and quantum mechanics, specifically the Schrödinger equation, where the operator evolving the observable is input independent.
@@ IV Summary and conclusions
Subsequently, we apply easy attention in the proposed transformer model for temporal-dynamics prediction of chaotic systems and we investigate the performance of our models on the Lorenz system. The transformer model with dense easy attention yields the lowest relative $l_{2}$-norm error of 2.0%, outperforming the same transformer architecture with the self attention (7.4%) and the LSTM model (37.7%). Furthermore, when inspecting the nine-equation model of wall-bounded turbulence by Moehlis et al. [42] the long-term statistics discussed in Fig. 5 show how LSTM yields very similar performance in the mean flow compared with the easy attention. Regarding the PDF of the Poincaré section shown in Fig. 5, the easy attention also exhibits better predictions than the other models, especially when considering the symmetries observed in Fig. 4 (top), which highlight the model’s ability to replicate invariant structures. Therefore, the easy-attention method emerges as the most accurate model when studying chaotic systems. Regarding the studied model of a nuclear reactor, the easy attention shows the best reconstructive capability when considering PSD and PDF analysis. The results also show that, compared with the self-attention-based transformer, the transformers using easy attention have a much lower computational cost, exhibit less complexity and have less parameters. Thus the transformer model using easy attention is more suitable for temporal-dynamics predictions.
Furthermore, a significant advantage of the easy-attention-based transformer is its accurate reconstruction of the invariant sets, as shown in Figs. 4, 6. These invariant sets arise when inspecting the systems independently of time, as the latter converge to a stationary probability density function and a fixed attractor in state space. The PDF is adequately reconstructed and the information density per neighborhood is accurately represented by the easy-attention model. In uniformly hyperbolic dynamical systems, the shadowing lemma guarantees that any learned or numerical pseudo-orbit given by a sequence $\{x_{k}\}$ satisfying:
$\|x_{k+1}-\Phi_{\Delta t}(x_{k})\|\leq\varepsilon\quad\text{for }k=0,\dots,N-1,$
with small $\varepsilon>0$, remains uniformly close to a true trajectory: there exists an exact orbit $\{\tilde{x}_{k}\}$ such that
$\|\tilde{x}_{k}-x_{k}\|\leq C\,\varepsilon\quad\text{for all }k,$
where $C>0$ depends on the system’s hyperbolicity constants [53, 54]. Hence, even when our transformer’s predictions diverge from the reference due to sensitive dependence on initial conditions, they form a valid pseudo-orbit that is shadowed by an actual orbit. This result justifies using the transformer’s output to recover statistically and geometrically meaningful features—such as invariant measures and attractor geometry—of the underlying chaotic system.
@@
V Methods
In this section, we will focus on the mathematical development of the attention part of the model and its implications on the final output, since the other blocks follow the standard transformer architecture [28]. To this end we first examine the self-attention model, then explain the easy-attention mechanism in detail. The mathematical implementation of a transformer encoder block is expressed as follows [55]: a transformer block is a parameterised function $T:\mathbb{R}^{n\times d}\rightarrow\mathbb{R}^{n\times d}$, usually consisting of an attention part and a feed forward-network, both with residual connections and normalization layer, see the transformer block used for each case in the Supplementary Material.